# Early Jamming Detection in Mobile Indoor Scenarios via Deep Learning

### v1

@@ V-B Experimental Settings
For our experiments, we found the best configuration of the hyperparameters of the autoencoder of BloodHound$+$, and we compared its performance against the approach based on the binary CNN Resnet-18 used in [7] and the solution based on 3-D images proposed in [8].
@@
Autoencoder deployment. We fine-tuned the hyperparameters of the autoencoder used in BloodHound$+$to find the best trade-off between classification accuracy and the general validity of the solution, i.e., to avoid overfitting. Specifically, we used the Matlab-provided implementation of autoencoders, version R2022b. As mentioned in Section IV, we used the Pure Linear (purelin) transfer function as the decoder transfer function. For all the remaining hyperparameters of the autoencoder, we used a formal hyperparameter optimization method conducted on a subset of our dataset. Specifically, we selected $1,500$ images from I-Q samples acquired in the scenario with the most data available, i.e., with the receiver positioned $10$ meters from the transmitter and the jammer emitting jamming with a RJP of $0.5$. We evaluated all combinations of the following hyperparameters: (i) hidden size (i.e., the size of the latent representation), with considered values being $8$, $16$, $32$ and $64$; (ii) sparsity regularization, with considered values being $1$, $0.5$, and 00; (iii) L2-regularization term, with considered values being $0.01$, $0.001$, and $0.0001$; and finally, (iv) encoder transfer function, with considered values being logsig and satlin. Finally, we set all remaining hyperparameters to their default values provided by Matlab, specified at [19]. Such a variety of configuration parameters led us to test a total number of $72$ configurations. For such tests, we used the methodology described below, inspired by the $k$-fold cross-validation technique used by the authors in [20].
@@
  • For each considered combination of hyperparameters $\alpha$, for every $i\in\{1,\dots,10\}$, we do the following:
@@
Note that, like the authors in [20], we selected the AUC as the optimization metric, since it measures the quality of a classifier independently of the threshold selection process. Following such a hyperparameter selection process, we set the hidden size value to $16$, the sparsity regularization term to $0.5$, the L2-regularization term to $0.01$ and the encoder transfer function to the Logistic Sigmoid (logsig). We also selected the number of epochs by examining how many iterations the autoencoder takes to reliably converge in the worst case, i.e., when setting the size of the latent representation to 64. We empirically established this to occur after 250 epochs, and thus, we selected such a value to trade off between classification accuracy and training time. We use such a hyperparameter configuration to train the autoencoder of BloodHound$+$ for all the results discussed below on unseen data, so as to avoid overfitting.
@@
Benchmark Approaches. For all experiments, we compare BloodHound$+$ with two benchmark solutions, i.e., the former version of BloodHound$+$ published in [7] and the approach proposed by the authors in [8]. We selected such solutions as they work successfully on PHY data, i.e., I-Q samples, significantly outperforming other solutions in challenging scenarios, such as the one considered in this manuscript. Both approaches use the residual CNN Resnet-18, pre-trained on the ImageNet dataset [21] and with the necessary modifications to the output layer necessary to fit the nature of the jamming detection problem. Specifically, the output layer of CNN has been modified to consider two possible classes as output, i.e., either No-Jamming or Jamming. Regarding the input layer, CNN works on images of size $224\times 224$ constructed over I-Q samples collected from the wireless channel, which is optimal for comparison to BloodHound$+$. To train CNN, we used the automated procedure trainNetwork provided by Matlab. We set the batch size to $32$, the number of epochs to $1$, and the solver to adam, similar to [7]. All the remaining hyperparameters have been set to their default values, available at [22].
@@
The solution provided in [7] works on grayscale images, while the proposal in [8] uses color images, characterized by three layers rather than one. To set up such images, we follow the same procedure as described in [8]. As an example, Fig. 6 shows the comparison of a grayscale and a color image generated over the same set of I-Q samples collected from the wireless channel.
@@
Figure 6: Sample grayscale and color unjammed (left) and jammed (right) images generated from the same set of I-Q samples, used for the experimental assessment. The jammed images have been obtained using samples collected at $10$ meters from the jammer, with RJP = $0.5$.
@@
Measurements Characterization. As an introduction to the presentation of our results, with reference to the setup using the USRP X310 SDRs, in Fig. 7 we show the BER of the TX-RX communication link experienced with different values of the RJP, under different jamming signals.
@@
Figure 7: Analysis of the BER of the TX-RX communication link using the SDRs USRP X310, with various levels of RJP and two jamming signals, i.e., tone jamming and Gaussian jamming. The receiver has been placed $10$ m away from the jammer.
@@
We highlight that, when $RJP<0.7$, the impact of the jammer on the communication link in terms of BER is very limited, i.e., very few bits are corrupted. On the contrary, when $RJP\geq 0.7$, most of the bits are corrupted. Recall that in this manuscript we are specifically interested in early jamming detection in a mobile scenario, i.e., detecting the jammer before its impact on the communication link becomes significant and significantly affects the throughput of the communication link. In this context, we are particularly interested in enhancing jamming detection performances when $RJP<0.7$. For higher values of $RJP$, other techniques based on the analysis of BER can already detect jamming ex-post, i.e., once the communication link is significantly affected.
@@
Finally, in this context, we note that the setup using the LimeSDR was configured with an absolute transmission gain at the jammer of $23$ dBm. This configuration allows us to have a BER of the legitimate communication link of approximately $0.001$, which allows us to study the impact of different configuration parameters of the scenario while matching the conditions of early jamming described above.
@@
Performance Metrics. We compare the performance of the methodologies introduced above mainly with reference to their accuracy, obtained as $acc=\frac{TP+TN}{TP+FP+FN+TN}$, being $TP$ the amount of true positives (i.e., jammed images correctly classified), $TN$ the true negatives (i.e., unjammed images correctly classified), $FP$ the false positives (i.e., unjammed images wrongly classified as jammed ones) and $FN$ the false negatives (i.e., jammed images wrongly classified as unjammed ones). For some of the results shown below, we show the True Positive Ratio (TPR) and True Negative Ratio (TNR), computed as $TPR=\frac{TP}{TP+FP}$ and $TNR=\frac{TN}{TN+FN}$, respectively. We obtain our estimates of accuracy, TPR, and TNR using the cross-validation approach used by the authors of [20], using 10-fold cross-validation. For each result, we report the mean and the $95\%$ confidence intervals, computed using the tool tinv provided by Matlab.
V-C Early Jamming Detection Robustness
@@
We first consider the impact of the received jamming power at the receiver on the capability of BloodHound$+$ to detect the presence of the jammer. We consider the setup using the Ettus X310 SDR, and specifically, the measurements where the receiver is located $10$ meters away from the jammer. Here, we generate images using $n=10^{5}$ I-Q samples and evaluate the performance of BloodHound$+$, the proposal in [8] and the solution in [7] to detect jamming, in terms of overall classification accuracy. We report the results of our investigation in Fig. 8.
@@
First, we note that the approaches in [8] and [7], based on CNNs, reliably identify jamming only when $RJP\geq 0.4$. With lower values of the RJP, their performances do not follow a unique trend, and the results also exhibit high variance (see the red bars indicating the confidence intervals of the measurements). On the contrary, BloodHound$+$reports remarkable performances for every value tested of $RJP$, with a minimum accuracy of $0.997$. We believe that such a result is due to the rationale of autoencoders used in BloodHound$+$, which work only on unjammed images. Such a configuration allows autoencoders to build a profile of the regular behavior of the wireless channel, in order to reliably identify smaller differences (anomalies).
We also investigated the impact of the distance of the receiver from the jammer on the performance of the cited solutions. To this aim, using the same setup as used for the previous tests, we set $RJP=0.5$ and move the receiver away from the jammer, up to a distance of $21$ meters. We stop at such a distance due to the physical limitations of the involved hardware, i.e., at distances higher than $21$ meters the BER of the legitimate communication link decreased significantly, preventing us from performing the test reliably. We report the results of our experiments in Fig. 9.
@@
Here, we notice that all the tested solutions report remarkable performance, with very limited differences. All three approaches reliably detect jamming at various distances, with an average accuracy of well above $0.99$. Overall, the distance between the receiver and the transmitters does not affect the classification accuracy, which remains very high even when the receiver is $21$ meters away.
Another important parameter of BloodHound$+$ is the number of samples used to generate images, namely $n$. The higher the value of such a parameter, the higher the number of samples to use for image generation, thus the longer the receiver has to acquire samples from the wireless channel, and the higher the processing overhead of the solution. For this analysis, in the same setup as in the previous experiments, we considered the data acquired with the receiver located $10$ m away from the jammer and $RJP=0.5$, and tested the performance of the three approaches while increasing the number of samples used in the image generation phase, from $10,000$ to $1,000,000$. Fig. 10 reports the results of our analysis.
@@
Note that the solutions in [8] and [7] report a higher classification accuracy than BloodHound$+$ for a low number of samples. For example, when the $n=10,000$, BloodHound$+$ reports an accuracy of $0.498$ while such values amount to $0.99$ and $0.95$ for the solution in [8] and [7], respectively. When the available number of samples increases, the accuracy of the benchmark approaches is still high, but the variance becomes larger (see the red bars). Conversely, when $n\geq 50,000$, not only the accuracy of BloodHound$+$ is very high (always higher than $0.99$), but the variance is also very limited (less than $0.001$ for all tests), which indicates greater robustness and reliability. Such results further motivate the deployment of BloodHound$+$ and highlight its superiority compared to the benchmark solutions.
To provide further insight into the performance of BloodHound$+$, we investigated the impact of additional configuration parameters. In particular, in the same setup as the last cited test, considering $n=100,000$, we evaluated the impact of the size of the training set. Figure 11 summarizes the results of our investigation, distinguishing the achieved TPR and TNR of our proposed solution.
@@
We notice that to perform reliably, BloodHound$+$requires a minimum training set of only $9$ images, reporting $TPR$ and $TNR$ values of $0.962$ and $0.99$, respectively. Performance stays almost constant when increasing the training set size. However, increasing the training set could be particularly relevant in very noisy scenarios, characterized by a wider range of expected wireless channel fluctuations. We recall that we need to train BloodHound$+$ only once, before deployment, and such results do not affect the deployability of our solution at runtime.
We also investigated further any bias of our results with respect to the specific hardware used for the experiments. Taking into account the same scenario as in previous experiments, i.e., the receiver located $10$ meters away from the jammer, we evaluated the TPR and TNR of BloodHound$+$ when changing the hardware used for jamming among the five available radios. We highlight that this methodology prevents the autoencoder from fingerprinting both the transmitter and the receiver, these being the same for all measurement classes. Moreover, we considered different hardware for the jammer during our measurements, i.e., we mutually excluded the ones adopted for training from the ones adopted for testing. This eventually guarantees that the autoencoder learns the characteristics of the legitimate signal only, while being independent of the transmitter, the receiver, and the jammer hardware (jammed signals are not used for training). Specifically, we consider all unjammed images obtained when placing the receiver at a distance of 10 meters from the transmitter, using radio $x$ as the jammer. Next, we considered all jammed images generated with the receiver located at a distance of 10 meters and using radio $y$ as a jammer. We separated the unjammed and jammed images into 10 folds, and we trained BloodHound$+$ on 9 of the folds containing unjammed images, holding the $i$-th one out to estimate the TNR. Next, we estimate the TNR on the $i$-th fold of unjammed images and estimate the TPR on the $i$-th fold of jammed images. Fig. 12 reports the result of our analysis. The tick labels on the x-axis are of the form $(x,y)$, as described above.
@@
Note that the TPR and TNR remain almost unchanged while varying the hardware used for jamming and the radio considered for training. Therefore, we can safely assume that BloodHound$+$ is not biased by the specific radios used in the experiments, but it can extract the features of the wireless channel useful for detecting jamming, being independent of specific hardware.
V-D Impact of Different Hardware and Sampling Rate
The results shown in the previous subsection have all been obtained using a single hardware brand (Ettus Research USRP X310) and a single configuration of the sample rate, i.e., $1$ Msa/s. To further extend our experimental assessment, we first ran further analyzes on the data collected through the second setup described in Section V-A, adopting the hardware LimeSDR Mini. An important consideration about the scenario described in Section III is that the jammer might not know in advance what the adopted sampling rate is. Thus, to disrupt ongoing communications as much as possible, in real-life scenarios, the jammer might emit jamming using the maximum achievable sampling ratio, likely higher than the one adopted by the legitimate communication link. At the same time, the receiver might oversample the signal, obtaining more useful information for jamming detection. This additional information might be discarded for communication purposes, but it might still be useful for jamming detection. Therefore, in our tests, we trained BloodHound$+$ on unjammed images of the legitimate communication link obtained under $RJP=0.5$ and a distance of 3 meters, and we tested using unjammed images (disjoint from the ones used for training) and jammed images obtained from I-Q samples generated with various levels of JOR. For this test, we let the jammer emit random Gaussian noise. We also compared the performance of BloodHound$+$ with the ones of the solution in [7]. Fig. 13 summarizes the results of our analysis, in terms of TPR (for comparison purposes).

### v4

@@ VI-B Experimental Settings
For our experiments, we found the best configuration of the hyperparameters of the autoencoder of BloodHound$+$ and compared its performance with the approach based on the binary CNN Resnet-18 used in [16] and the solution based on 3-D images proposed in [17].
@@
Autoencoder deployment. We fine-tuned the hyperparameters of the autoencoder used in BloodHound$+$to find the best trade-off between classification accuracy and the general validity of the solution, i.e., to avoid overfitting. Specifically, we used the Matlab-provided implementation of autoencoders, version R2022b. As mentioned in Section V, we used the Pure Linear (purelin) transfer function as the decoder transfer function. We used a formal hyperparameter optimization method performed on a subset of our dataset for all remaining hyperparameters of the autoencoder. Specifically, we selected $1,500$ images from I-Q samples acquired in the scenario with the most data available, i.e., with the receiver positioned $10$ meters from the transmitter and the jammer emitting jamming with a RJP of $0.5$. We evaluated all combinations of the following hyperparameters: (i) hidden size (i.e., the size of the latent representation), with considered values being $8$, $16$, $32$ and $64$; (ii) sparsity regularization, with considered values being $1$, $0.5$, and 00; (iii) L2-regularization term, with considered values being $0.01$, $0.001$, and $0.0001$; and finally, (iv) encoder transfer function, with considered values being logsig and satlin. Finally, we set all remaining hyperparameters to their default values provided by Matlab, specified at [53]. Such various configuration parameters led us to test a total number of $72$ configurations. For such tests, we used the methodology described in the following, inspired by the $k$-fold cross-validation technique used by the authors in [54].
@@
  • For each combination of hyperparameters considered $\alpha$, for every $i\in\{1,\dots,10\}$, we do the following:
@@
Note that, like the authors in [54], we selected the AUC as the optimization metric since it measures the quality of a classifier independently of the threshold selection process. Following such a hyperparameter selection process, we set the hidden size value to $16$, the sparsity regularization term to $0.5$, the L2-regularization term to $0.01$ and the encoder transfer function to the Logistic Sigmoid (logsig). We also selected the number of epochs by examining how many iterations the autoencoder takes to reliably converge in the worst case, i.e., when setting the size of the latent representation to 64. We empirically established that this would occur after 250 epochs, and thus, we selected such a value to trade-off between classification accuracy and training time. We use such a hyperparameter configuration to train the autoencoder of BloodHound$+$ for all the results discussed below on unseen data to avoid overfitting.
@@
Benchmark Approaches. For all experiments, we compare BloodHound$+$ with two benchmark solutions, i.e., the former version of BloodHound$+$ published in [16] and the approach proposed by the authors in [17]. We selected such solutions as they work successfully on PHY data, i.e., I-Q samples, significantly outperforming other solutions in challenging scenarios, such as the one considered in this manuscript. Both approaches use the residual CNN Resnet-18, pre-trained on the ImageNet dataset [55] and with the necessary modifications to the output layer necessary to fit the nature of the jamming detection problem. Specifically, we modified the output layer of CNN to consider two possible classes as output, i.e., either No-Jamming or Jamming. Regarding the input layer, CNN works on images of size $224\times 224$ constructed over I-Q samples collected from the wireless channel, which is optimal for comparison to BloodHound$+$. To train CNN, we used the automated procedure trainNetwork provided by Matlab. We set the batch size to $32$, the number of epochs to $1$, and the solver to adam, similar to [16]. We set all remaining hyperparameters to their default values, available at [56].
@@
The solution provided in [16] works on grayscale images, while the proposal in [17] uses color images characterized by three layers rather than one. We follow the same procedure described in [17] to set up such images. As an example, Fig. 6 compares a grayscale and a color image generated over the same set of I-Q samples collected from the wireless channel. The lower images in Fig. 6 have been generated by considering one layer for each primary colour component (red, green, and blue). Therefore, assuming an image constituted by a three-layer matrix, i.e., [224 × 224 × 3] (one layer for each primary colour), and the pixel value between 0 and 255, in line with [17], we assign each value of the tile through the following rule.
  • $0\leq x_{T}\leq 255$, then $p_{R}=0,p_{G}=0,p_{B}=x_{T}$,
  • $256\leq x_{T}\leq 511$, then $p_{R}=0,p_{G}=x_{T}-255,p_{B}=255$,
  • $x_{T}>511$, then , then $p_{R}=x_{T}-510,p_{G}=255,p_{B}=255$,
where $x_{T}$ represents the value of the tile from the bi-variate histogram, while $p_{R},p_{G}$ and $p_{B}$ are the pixel values, i.e., red, green and blue, respectively. Finally, we observe that if $x_{T}>767$, it is clipped to 767—this issue can also be controlled by properly adjusting the chunk size. Instead, for the upper figures, in line with the logic in [16], we only consider one layer and thus, $0\leq x_{T}\leq 255$.
@@
Figure 6: Sample grayscale and colour unjammed (left) and jammed (right) images generated from the same set of I-Q samples, used for the experimental assessment. We generated the jammed images using samples collected at $10$ meters from the jammer, with RJP = $0.5$.
@@
Measurements Characterization. As an introduction to the presentation of our results, with reference to the setup using the USRP X310 SDRs, in Fig. 7, we show the BER of the TX-RX communication link experienced with different values of the RJP, under different jamming signals.
@@
Figure 7: Analysis of the BER of the TX-RX communication link using the SDRs USRP X310, with various levels of RJP and two jamming signals, i.e., tone jamming and Gaussian jamming. We placed the receiver $10$ m away from the jammer.
@@
We highlight that when $RJP<0.7$, the impact of the jammer on the communication link in terms of BER is minimal, i.e., very few bits are corrupted. We achieved the lowest BER value with $RJP=0.1$, where $BER\approx 1e-6$. On the contrary, most bits are corrupted when $RJP\geq 0.7$. Recall that in this manuscript, we are specifically interested in detecting jamming in a mobile scenario in a low-BER regime, i.e., before its impact on the communication link becomes significant and significantly affects the throughput of the communication link. In this context, we are particularly interested in improving the jamming detection performance when $RJP<0.7$. For higher values of $RJP$, other techniques based on the analysis of BER can already detect jamming ex-post, i.e., once the communication link is significantly affected.
@@
Finally, in this context, note that we configured the LimeSDR setup with an absolute transmission gain at the jammer of $23$ dBm. This configuration allows us to have a BER of the legitimate communication link of approximately $0.001$, which enables us to study the impact of different configuration parameters of the scenario while matching the conditions of the low-BER regime described above.
@@
Performance Metrics. We compare the performance of the methodologies introduced above mainly concerning their accuracy, obtained as $acc=\frac{TP+TN}{TP+FP+FN+TN}$, being $TP$ the number of true positives (i.e., jammed images correctly classified), $TN$ the true negatives (i.e., unjammed images correctly classified), $FP$ the false positives (i.e., unjammed images wrongly classified as jammed ones) and $FN$ the false negatives (i.e., jammed images wrongly classified as unjammed ones). For some of the results shown below, we show the True Positive Ratio (TPR) and True Negative Ratio (TNR), computed as $TPR=\frac{TP}{TP+FP}$ and $TNR=\frac{TN}{TN+FN}$, respectively. We obtain our estimates of accuracy, TPR, and TNR using the cross-validation approach used by the authors of [54], using 10-fold cross-validation. For each result, we report the mean and the $95\%$ confidence intervals, computed using the tool tinv provided by Matlab. Tab. III summarizes all the experiments discussed below, together with the relevant parameters.
TABLE III: Experiments carried out in the manuscript and related parameters tested.
Techniques RJP Distances [m] Samples per Image Training Set Size Jamming Radios Jamming Oversampling Rates Jamming Signal Type
BloodHound$+$
0.1 - 0.8
3, 5, 7, 10, 13, 16, 19, 21
10,000 50,000 100,000 500,000 1,000,000
2, 9, 18, 36, 54, 72
4, 5, 6, 7
1, 2, 3, 4
AWGN, BPSK
[16]
0.1 - 0.8
3, 5, 7, 10, 13, 16, 19, 21
10,000 50,000 100,000 500,000 1,000,000
1, 2, 3, 4
AWGN, BPSK
[17]
0.1 - 0.8
3, 5, 7, 10, 13, 16, 19, 21
10,000 50,000 100,000 500,000 1,000,000
VI-C Jamming Detection Robustness in Low-BER Regime
@@
We first consider the impact of the received jamming power at the receiver on the capability of BloodHound$+$ to detect the presence of the jammer. We consider the setup using the Ettus X310 SDR, and precisely, the measurements where we placed the receiver $10$ meters away from the jammer. Here, we generate images using $n=10^{5}$ I-Q samples and evaluate the performance of BloodHound$+$, the proposal in [17] and the solution in [16] to detect jamming, in terms of overall classification accuracy. We report the results of our investigation in Fig. 8.
@@
First, we note that the approaches in [17] and [16], based on CNNs, reliably identify jamming only when $RJP\geq 0.4$. With lower values of the RJP, their performances do not follow a unique trend, and the results also exhibit high variance (see the red bars indicating the confidence intervals of the measurements). On the contrary, BloodHound$+$ reports remarkable performances for every value tested of $RJP$, with a minimum accuracy of $0.997$. We believe such a result is due to the rationale of autoencoders used in BloodHound$+$, which work only on unjammed images. This configuration allows auto-encoders to build a profile of the regular behavior of the wireless channel to identify more minor differences (anomalies) reliably. We also note that the performances of BloodHound$+$ do not depend on the speed of the involved entities nor the smoothness of the change of channel conditions. Indeed, our approach processes chunks of $n$ I-Q samples and compares the images created from such samples with the expected channel conditions acquired during training time.
We also investigated the impact of the receiver’s distance from the jammer on the performance of the cited solutions. To this end, using the same setup used for the previous tests, we set $RJP=0.5$ and move the receiver away from the jammer to a distance of $21$ meters. We stopped at such a distance due to the physical limitations of the involved hardware, i.e., at distances higher than $21$ meters, the BER of the legitimate communication link increases significantly, preventing us from performing the test reliably. We first acknowledge that the receiver experiences increasing interference while moving closer to the jammer. This phenomenon is shown in Fig. 9, highlighting that the process is not linear but follows a polynomial model while increasing distance, i.e., $ax^{2}+bx+c$. Note that our result is consistent with the findings of Tedeschi et al. reported in [57].
Figure 9: Normalized Received Signal Strength (NRSS) as a function of the distance: the NRSS decreases when the receiver moves far away from the transmitter. The solid red line shows the model that best fits the experimental data (black circles).
We also notice that, for specific distance values, significantly different RSS values may be experienced, due to the fast-fading process affecting wireless communication channels. However, recall that lower RSS does not necessarily imply higher BER. As our manuscript focuses on jamming detection in low-BER scenarios, our analysis considers a condition where the BER of the communication channel is low to provide jamming detection before the loss of the communication link. We report in Fig. 10 the results of our experiments involving jamming detection via BloodHound$+$ and competing approaches.
@@
Here, we notice that all tested solutions report remarkable performance with minimal differences. All three approaches reliably detect jamming at various distances, with an average accuracy well above $0.99$. Overall, the distance between the receiver and the transmitters does not affect the classification accuracy, which remains very high even when the receiver is $21$ meters away. We present such a result primarily to show that, similarly to competing approaches, BloodHound$+$ can detect jamming even at a significant distance from the jamming source. In addition to such a result, we present additional results below that show the enhanced robustness of BloodHound$+$ compared to other approaches.
Another critical parameter of BloodHound$+$ is the number of samples used to generate images, namely $n$. The higher the value of this parameter, the higher the number of samples to use for image generation. Thus, the longer the receiver has to acquire samples from the wireless channel, the higher the processing overhead of the solution. For this analysis, in the same setup as in the previous experiments, we considered the data acquired with the receiver located $10$ m away from the jammer and $RJP=0.5$, and tested the performance of the three approaches while increasing the number of samples used in the image generation phase, from $10,000$ to $1,000,000$. Fig. 11 reports the results of our analysis.
@@
Note that the solutions in [17] and [16] report a higher classification accuracy than BloodHound$+$ for a low number of samples. For example, when the $n=10,000$, BloodHound$+$ reports an accuracy of $0.498$ while such values amount to $0.99$ and $0.95$ for the solution in [17] and [16], respectively. When the available number of samples increases, the accuracy of the benchmark approaches is still high, but the variance becomes larger (see the red bars). On the contrary, when $n\geq 50,000$, not only the accuracy of BloodHound$+$ is very high (always higher than $0.99$), but the variance is also minimal (less than $0.001$ for all tests), indicating greater robustness and reliability. Such results further motivate the deployment of BloodHound$+$ and highlight its superiority compared to the benchmark solutions.
To provide further insight into the performance of BloodHound$+$, we investigated the impact of additional configuration parameters. In particular, in the same setup as the last cited test, considering $n=100,000$, we evaluated the effect of the training set size. Figure 12 summarizes the results of our investigation, distinguishing the achieved TPR and TNR of our proposed solution.
@@
To perform reliably, BloodHound$+$ requires a minimum training set of only $9$ images, reporting $TPR$ and $TNR$ values of $0.962$ and $0.99$, respectively. Performance remains almost constant when increasing the training set size. However, increasing the training set could be particularly relevant in very noisy scenarios characterized by a broader range of expected wireless channel fluctuations. We recall that we need to train BloodHound$+$ only once before deployment, and such results do not affect the deployability of our solution at runtime.
We also investigated further any bias of our results concerning the specific hardware used for the experiments. Taking into account the same scenario as in previous experiments, i.e., the receiver located $10$ meters away from the jammer, we evaluated the TPR and TNR of BloodHound$+$ when changing the hardware used for jamming among the five available radios. We highlight that this methodology prevents the autoencoder from fingerprinting both the transmitter and the receiver, these being the same for all measurement classes. Moreover, we considered different hardware for the jammer during our measurements, i.e., we mutually excluded the ones adopted for training from the ones adopted for testing. The mentioned strategy eventually guarantees that the autoencoder learns the characteristics of the legitimate signal only while being independent of the transmitter, the receiver, and the jammer hardware (we do not use jammed signals for training). Specifically, we consider all unjammed images obtained when placing the receiver at a distance of 10 meters from the transmitter, using the radio $x$ as the jammer. Next, we consider all the jammed images generated with the receiver located at a distance of 10 meters and use radio $y$ as a jammer. We separated the unjammed and jammed images into 10 folds, and we trained BloodHound$+$ on 9 of the folds containing unjammed images, holding the $i$-th one to estimate the TNR. Next, we estimate the TNR on the $i$-th fold of the unjammed images and evaluate the TPR on the $i$-th fold of the jammed images. Fig. 13 reports the result of our analysis. The tick labels on the x-axis are of the form $(x,y)$, as described above.
@@
Note that the TPR and TNR remain almost unchanged while varying the hardware used for jamming and the radio considered for training. Therefore, we can safely assume that BloodHound$+$ is not biased by the specific radios used in the experiments. Still, it can extract the features of the wireless channel useful for detecting jamming independently of the particular hardware.
VI-D Impact of Different Hardware and Sampling Rate
We obtained all the results shown in the previous subsection using a single hardware brand (Ettus Research USRP X310) and a single configuration of the sample rate, i.e., $1$ Msa/s. To further extend our experimental assessment, we first analyzed the data collected through the second setup described in Section VI-A, adopting the hardware LimeSDR Mini. An important consideration about the scenario described in Section IV is that the jammer might not know the adopted sampling rate in advance. Thus, to disrupt ongoing communications as much as possible, in real-life scenarios, the jammer might emit jamming using the maximum achievable sampling ratio, likely higher than the one adopted by the legitimate communication link. At the same time, the receiver might oversample the signal, obtaining more helpful information for jamming detection. This additional information might be discarded for communication but might benefit jamming detection. Therefore, in our tests, we trained BloodHound$+$ on unjammed images of the legitimate communication link obtained under $RJP=0.5$ and a distance of 3 meters. Then, we tested using unjammed images (disjoint from those used for training) and jammed images obtained from I-Q samples generated with various levels of JOR. For this test, we let the jammer emit random Gaussian noise. We also compared the performance of BloodHound$+$ with the solution in [16]. Fig. 14 summarizes the results of our analysis in terms of TPR (for comparison purposes).